A car A and B travels towards each other on a 42-km straight highway. Car A travels at a rate of 60 kph while car B travels at a rate of 80 kph. In how many minutes will the 2 care meet?
a.6 min
b.18 min
c.30 min
d.42 min
e.84 min

Answer by ankor@dixie-net.com(22740)

A car A and B travels towards each other on a 42-km straight highway.
Car A travels at a rate of 60 kph while car B travels at a rate of 80 kph.
In how many minutes will the 2 cars meet?
:
When the two cars meet, they will have traveled a total 42 km
let t = time traveled when they meet
write a dist equation; dist = speed * time
60t + 80t = 42
140t = 42
t = 42/140
t = .3 hrs
Change to min
.3 * 60 = 18 min